Thomas Andrews (Wilfred Shadbolt) is somewhat concerned that he keeps being cast as unpleasant persons. He is known for his sweet and even temperament. You might have seen him playing Pooh-Bah in Mikado or Dick Deadeye in HMS Pinafore.

Jeffrey Bellin (Music Director) While Jeffrey daylights for Dow Jones in the Wall Street Journal offices, he comes out of his corporate cocoon to revel in his professional career interests of acting and playing music. In addition to his background as a violist and his current work as the Music Director, Jeffrey is also studying improv with Boston’s leading troupe, Improv Boston. Previous credits include A Funny Thing Happened on the Way to the Forum and $30 worth of merchandise at T. J. Maxx.

Gillian P. Bidgood W’93 (Helen Smith, Citizen) After serving time on the Executive Committee, Jill is happy to have snatched her first singing appearance with the MIT G&SP.

Felicia Brady (C) (Kate Carruthers) Contrary to numerous rumors, Felicia is not Cindy’s illegitimate younger sister sold to the black market. For some unknown, possibly masochistic, reason, she has chosen to get her theatrical feet wet for the first time doing G&S. In real life, she is a pianist. In fact, if you are considering taking private lessons, call her at 864-4104.

Ginny Briggs (Elsie Maynard) has played soprano roles in other G&S shows including Iolanthe, Patience, and most recently as Josephine in HMS Pinafore with the Concord Players. She has performed with various community theatres in Dames At Sea, Fiddler on the Roof and the title role in Gigi. She’d like to say she misses the Kahlua Hawaiian and Flapball-changes with her “travelling” group.

Mike Bromberg ’70 (Lighting Designer) is in Lighting Heaven. After five shows in 54-100, he appreciates twice the light for half the price, not to mention a lock on the door. When not lighting G&S, he climbs mountains, plays the piano, and draws maps.

Thomas Burhoe Chadwick, Jr. ’93 (Colonel Fairfax) Although this is his first role with MIT G&SP, Thomas is no stranger to the stage. For the past seven years, Thomas has worked with, quite frankly, POoP (Popular Opera of Pittsfield—A littleknown company from his hometown in Maine), where he has had the opportunity to rub elbows with performers from New York G&SP. After playing a bearded, blind pirate, and a space-age apeman, he is relieved to find this role requires additional face hair for only one scene.

Richard Davis ’94 (Dick Smith, Citizen) This show is his second at MIT, the first being this past summer’s performance of Sweeney Todd by MTG where he appeared as Pirelli. In other activities, he heads the MIT chapter of the Flat Earth Society, and his highest aspiration is to become a class 5 free roaming vapor. He is also in course 6. AAAAAH!

Alan deLespinasse ’94 (Patrick Ernest Walker, Jr., Citizen) Are you bored out of your cranium? Sick of listening to mice philosophizing on the radio? Tired of reading in circles? Come to the only really happening place in town—in fact, in this entire multidimensional chromosome! It’s The Furniture Department! With disgustingly live shows and highly chemically reactive food nearly every night! Tonight, due to poor organization, the Paranoid Gerbils will not be appearing as previously announced. Instead, we present a medieval cyclic play with an unbroadcastable title! Come and have a bouncy rubber sphere wasting your money and intoxicating your mind! Remember, The Furniture Department is a Fun Place! [Stupid commercial theme.]

John Fesenko (G) (Jonathan Farmer, Yeoman) is a masochistic graduate student (forgive the redundancy) who, determined to snuff out of existence any scant free time which he still had, joined this group two weeks into rehearsals, ensuring a month chock full of glassy-eyed all-nighters and unresolved existential dilemmas. In fact, the experience has been so satisfying in this regard that he will no doubt appear again and again over the course of his five-year hiatus at MIT.

Mary A. Finn ’81 (Costume Designer; Margaret Brewster, Citizen) looked at us with anguished eyes and whimpered, “But it’s hard to write a bio when you haven’t got a life!”

Erik D. Harder (C) (Sergeant Meryll) This is Erik’s first G&S appearance since his debut in a fourth-grade production of H.M.S. Pinafore nearly two decades ago. He hopes it won’t be quite as long ’till the next one. By day, Erik pretends to work with computers at the Whitehead Institute. By night, he dozes through linguistics classes at Harvard.

Zoot F. Harrison (C) (Jack Point) “Zoot, Zoot, Zoot,” said a friend of mine, “Why do you do theater?” Zoot would like to thank Luiz, El Gallo, Florian, Pish Tush, Rusty Charlie, Colonel Calverly, Sir Roderick, Sir Marmaduke Poindexter, Lord Mountararat, Captain Corcoran, Ko-Ko, and Judge Turpin: dudes Zoot has known. As Dave’s alterego, Zoot continues to ruin his weal-crafted, linear, goal-oriented career plans by being silly.

Derek Herrera ’92 (Leonard Meryll; The Headsman) will someday rule the entire universe!!! Nothing shall stand in his way of total, complete domination of all forms of life!!! Succumb all to the might of plaid!!! HA HA HA HA!!!! (Note: Coca-Cola is very, very bad for your mental stability).

Michael

Hoch ’90 (Lieutenant, Sir Richard Cholmondely) In his farewell performance in the Boston area, Michael enjoys actually speaking on stage and fears an inexcusable lack of G&S in Farmtown, Maryland, U.S.A.

Andrea Humez (Sophie, Citizen) Still the youngest member of the cast, (she’ll be in college next year. Really!), Andrea defies typecasting by playing her third fifteenyear-old chorister. She is still not quite sure what she’s doing here, but if her name appears in the program next year, it will mean there must have been some reason. She is also not sure why she is singing soprano.

Kevin Iga ’92 (James Nasium (Citizen); Programme) As a member of too many student organizations, Kevin wonders why he is singing in this show. He also wonders why he is singing tenor in this show. But, then again, he also wonders why H_{deR}^{p}(M) ≅ ∞H_{p}(M; ℜ)^{∗}, so the fact that he’s double majoring is not at all disturbing to him. He would like to remind those reading this in the first weekend performance to give blood at the Red Cross TCA blood drive in Kresge Lobby, and to vote at the APO “Ugliest Man On Campus” competition charity fundraiser in Lobby 10.

Frank Jones (C)

(George Robertson, Yeoman; Master Carpenter; Assistant Producer) has appeared as chorus member in one other MIT G&SP show called Crimes against Japanese Culture, or The Mikado. He is functioning as Jackof-all-trades in this production: Yeoman, Master Carpenter, and Assistant Producer. In real life, he is a bookbinder working for MIT and a student at Longy School studying early Double Reeds.

Kathleen Keegan (Phoebe Meryll) Yeoman is Kathy’s second production with MIT G&SP. Last spring she played a singing schoolgirl in the Mikado. Kathy enjoys acting, singing and improv and would love to make a career of all three, yet there are those who remain convinced that she should pursue her lifelong dream to sell Tupperware^{T M}.

Deborah Kreuze ’91 (Dame Carruthers, Assistant Director) would like those members of the audience who care to to ponder briefly the reality of the institutionalized torture and killing upon which this show is based, and then to ask themselves why Gilbert would choose to have an unmarried, middle-aged woman sing the song extolling the virtues of such violence. (25 points; partial credit will be given.)

Adam Lindsay ’94 (Andrew Stewart, 1st Yeoman) has been plum tickled to appear in his seventh show here at Tufts as Julie Jordan. This is the least ditzy soprano he’s played to date, and he really enjoyed being a mother on stage. All in all, this show has been a “real nice clambake”. He’d like to give everyone in the cast, crew, and orchestra a big wet smooch and a hug, and remind them that “You’ll never walk alone...”

Karen Mueller (C) (Stage Director) directed this spring’s MIT G&SP production of Mikado, and has also served with the group as Tech Director, Set Designer, Set Crew, and Rapturous Maiden. She has directed G&S at Oberlin College, worked as ASM at the New Rep Theatre, and been involved in various other theatrical endeavors for 13 years. By day, she devotes herself to Excellence in Public Service at MIT’s Rotch Library.

Joan Shear (Choreographer) “Smiley the Choreographer” also choreographed the 1983 Harvard Law School Show, “North by North Middle,” as well as many dances for her sister and herself when they were children. More recently she frequently teaches folk dances at the MIT Folk Dance Club.

Matt Sikorski (C) (Scenic Artist, a.k.a. Set Designer) designed the set for MIT G&SP’s The Mikado and would like to reside somewhere between the pure gesture and the expressive form.

CarolYN Jean Smith ’87 (Sarah the Barmaid, Citizen) has been a part of this organization since its inception in 1988, and after performing in eight shows (seven as soprano chorus, one as court musician), producing five shows, house managing one show and being on Exec-Comm for two years she is very happy to be doing chorus and costumes, period. And she’s very glad that her hair is no longer green.

Mr.

Spoon (Jack’s Bauble) This is Mr. Spoon’s first public performance. He joins the MIT G&SP after a successful career baking cookies (where he made a lot of dough). Originally just a rehearsal prop, he decided to get into the limelight by imitating his hero, Milton Berle, who also dressed up in women’s underwear.

J.B. Sweeney (C) (Jane Alexaneder, Citizen; Publicity) is often found playing guitar and singing folk songs around campfires, though her secret passion is playing drum set with rock bands. (It’s not such a secret anymore...) This is J.B.’s first experience singing music that wasn’t written during this century, and she could see herself making a habit of it. Rapture! Rapture!

���� ��˘������ (Yanko Sheiretov) ’92 (Jean-Claude Manier, 2.5^{th} Yeoman) If all goes well, this is Yanko’s last appearance in MIT G&SP as an MIT undergraduate. He was previously seen in Sorcerer, Iolanthe, Pinafore, and Mikado. Being a senior now with a workload of only two classes and a thesis, he is happy that schoolwork can no longer impede his brilliant musical theatre career. He has strong faith in the Second Law of Thermodynamics.

Laura White (Producer; Properties Manager; Co-Publicity Manager; Counsellor; Occasional Vocal Coach; MIT G&SP member and frequent leading player since 1988) I see clearly and with joy what I have become and what I will be. Therefore, I must leave this company. Goodbye, MIT G&SP. Thank you for everything.